OpenStudy (anonymous):

what is the 100th derivative of f(x) = sinx ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

sin(x), no joke

OpenStudy (anonymous):

I also think it is sin(x)

OpenStudy (anonymous):

it's even ... should be cos

OpenStudy (anonymous):

cause 1 derivative would be cos, then -sin, then -cos, then sin again, which would be the 4 derivative and 100 is multiple of 4

OpenStudy (anonymous):

it's not cos, it goes: sin(x) cos(x) -sin(x) -cos(x) every fourth one is the same, hence 100 would be the 25th "fourth one"

OpenStudy (anonymous):

agreed "-cos" :) cool!

OpenStudy (anonymous):

it is sin(x) not -cos

OpenStudy (anonymous):

how come? see dirtbikr remarks...

OpenStudy (anonymous):

sin is not the 1 derivative of sin, cos is the 1 one

OpenStudy (anonymous):

dude... you are right! sin (x) !

OpenStudy (anonymous):

f'(x)=cosx=sin(pi/2+x) f''(x)= -sin(x)= sin(pi/2+x) f'''= -cosx= sin(3pi/2+x) f''''= sinx= sin(4pi/2+x) f(100)= sin(100pi/2+x)=sinx